OpenStudy (anonymous):

I need help someone please explain Ok so k12 gave an example: 1/8 + 3/5 1/8 + 3/5= 5/40 + 24/40 =29/40 So, my question is, how did they get 5 from 5/40? I know how they got 40, all you have to do is find the LCD ,but how did they get the numerator, 5? Idk maybe it's probably obvious ,but I don't get it, maybe i'm missing something from the problem??

Mehek (mehek14):

when adding fractions, you need the same denominators so the common denominator is 40 1/8 = 5/40 because 8 * 5 = 40 1 * 5 = 5 5/40

OpenStudy (anonymous):

yeah that is correct since u did 8 times 5 to get 40, u do the same thing 4 the top 5 time 1

OpenStudy (anonymous):

Thanks!